OpenStudy (anonymous):

(−5 + 3i) • (1 − 2i) PLEASE someone explain this to me?

OpenStudy (sweetburger):

is i considered \[\sqrt{-1}\]

OpenStudy (anonymous):

yes, i'm supposed to simplify

OpenStudy (math&ing001):

Did you distribute (a+b)(c+d)=ac+ad+bc+bd If yes tell me where you're stuck.

OpenStudy (sweetburger):

foil method

OpenStudy (math&ing001):

(-5+3i)(1-2i) =-5+10i+3i-6i*i i*i=-1 =-5+13i+6 =1+13i